An elevator cage is going up with a velocity of 6 m/s. When the cage was 36 m above the bottom of theshaft, a bolt gets detached from the bottom of the cage floor. Find the velocity with which the bolt strikesthe bottom of the shaft and the time that elapses.


1
Expert's answer
2021-05-27T17:48:32-0400

"h=\frac{gt^2}{2},\implies t=\sqrt{\frac{2h}{g}}=2.7~s,"

"v=gt=\sqrt{2gh}=26.6~\frac ms."